Solve the following initial value problems using method of undetermined coefficients: (i) y−9y = ex+x−1, y(0) = −1,y(0) = 1,


1
Expert's answer
2022-02-28T12:07:48-0500

Correponding homogeneous differential equation


"y''-9y=0"

Characteristic (auxiliary) equation


"r^2-9=0"

"r_1=3, r_2=-3"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{3x}+c_2e^{-3x}"

Find the particular solution of the non homogeneous differential equation


"y_p=Ae^x+Bx+C"

"y_p'=Ae^x+B"

"y_p''=Ae^x"

Substitute


"Ae^x-9Ae^x-9Bx-9C=e^x+x-1"

"A=-1\/8"

"B=-1\/9"

"C=1\/9"

The general solution of the non homogeneous differential equation is


"y=c_1e^{3x}+c_2e^{-3x}-\dfrac{1}{8}e^x-\dfrac{1}{9}x+\dfrac{1}{9}"

"y'(0) = \u22121,y(0) = 1"


"3c_1-3c_2-\dfrac{1}{8}-\dfrac{1}{9}=-1"

"c_1+c_2-\dfrac{1}{8}+\dfrac{1}{9}=1"

"c_2=\dfrac{73}{72}-c_1"

"6c_1-\dfrac{219}{72}=-\dfrac{55}{72}"

"c_1=\dfrac{82}{216}"


"c_2=\dfrac{137}{216}"

The solution of the given IVP is


"y=\dfrac{82}{216}e^{3x}+\dfrac{137}{216}e^{-3x}-\dfrac{1}{8}e^x-\dfrac{1}{9}x+\dfrac{1}{9}"